U=1,2,3,4,5,6,7
A=1,2,4,5
B=1,3,5,7
How Many Elements Are In A U B'

Found 3 solutions by Edwin McCravy, solver91311, ikleyn:
Answer by Edwin McCravy(20055)


When you list the elements in a set always put braces {   }
around them.  I noticed you did not do that and I put them
for you, but you should always put braces around a list of
elements. 

A U B'

Substitute {1,2,4,5} for A

{1,2,4,5} U B'

Substitute {1,3,5,7} for B

{1,2,4,5} U {1,3,5,7}'

The prime ' means the complement set which is all the elements
in the universal set besides those which are in B, which are

So we replace {1,3,5,7}' by {2,4,6}

{1,2,4,5} U {2,4,6}

Union means we combine those two sets into one:

{1,2,4,5,6}    <--Answer  (We do not list 2 and 4 but once)

Edwin


Answer by solver91311(24713)







John

My calculator said it, I believe it, that settles it


Answer by ikleyn(52781)

.

    Use the formula


        n(A U B') = n(A) + n(B') - n(A ∩ B').


     In the formula


         n(A) = 4;  n(B') = 3;  n(A ∩ B') = 2.


     Hence,  n(A U B') = 4 + 3 - 2 = 5.      ANSWER

---------

Be aware :   in the post by  John,  there is a small deficiency,  which makes his formula incorrect.

It is why I wrote my post.